STATEMENT OF COMPREHENSIVE INCOME 

a. COGS (Cost of Good Sold)

b. NET INCOME

C. COMPREHENSIVE INCOME


1. The Swift Company presented the balances of each income and expense account and you are tasked to make its expanded income statement for the year ended 2020.

 2. Always apply the income tax expense as 25% of operating profit. 


Gross Sales = ₱ 3,822,300 

Sales Returns= ₱ 219,700 

Gross Purchases = ₱ 1,000,000 

Purchase Discounts = ₱ 180,000 

Beginning Inventories = ₱ 920,500

 Ending Inventories = ₱ 450,800

 Distribution Expenses = ₱ 760,200

 Administrative Expenses = ₱ 440,900 

Other Comprehensive Income = ₱ 210,000



1
Expert's answer
2021-10-07T09:40:38-0400

a. COGS (Cost of Good Sold)

= Beginning Inventories + Purchases -Ending Inventories

Net Purchases = Gross Purchases - Purchases Discounts

= 1,000,000 - 180,000 = 820,000

COGS = 920,500 + 820,000 - 450,800 = 1,289,700


b. NET INCOME

Net Income = Gross Profit - all other Expenses


Gross Profit = Net sales - COGS

Net Sales = Gross Sales - Sales Returns

= 3,822,300 - 219,700 = 3,602,600

Gross Profit = 3,602,600 - 1,289,700 = 2,312,900


Operating Profit = Gross Profit - Operating Expenses

= Gross Profit - (Distribution Expenses + Administrative Expenses)

= 2,312,900 - 760200 + 440,900 = 1,111,800


Income Tax Expense = 25% × 1,111,800 = 277,950


Net Income = Gross Profit -(Distribution Expenses + Administrative Expenses + Income Tax Expense)

= 2,312,900 - 1,479,050 = 833,850


C. COMPREHENSIVE INCOME

Swift Company

Statement of Comprehensive Income

For the Year ended 2020


Sales. 3,602,600

COGS. ( 1,289,700)

Gross Profit. 2,312,900

Operating Expenses

Distribution Expenses. 760,200

Administrative Expenses. 440,900(1,201,100)

Income from continuing operations before income tax. 1,111,800

Income tax Expense. 277,950

Income from continuing operations. 833,850

Net Income for the year ended 2020. 833,850